# Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition

### v2

@@ 4.4 Error Patterns and Analysis
Our analysis reveals two primary patterns in the overreaction problem (false positives), where models misclassify safe situations as dangerous, as illustrated in Figure 4:
@@
Our analysis across 14 models revealed that Contextual Overinterpretation was dominant, accounting for 88-93% of all misclassifications regardless of model architecture or scale. Notably, in the Natural Disasters category, 100% of the errors were Contextual Overinterpretation, suggesting that models can correctly identify elements like fire or water but consistently fail to assess their contextual safety.
These findings suggest that current VLMs can detect potentially hazardous elements, but lack the nuanced reasoning required to assess whether these elements pose actual dangers in specific contexts—a critical limitation for their reliability in safety applications. Detailed error pattern analysis across model sizes and categories, along with additional examples of Contextual Overinterpretation, are provided in the supplementary material.
@@ 4.5 Model Size and Category Effects
Our analysis reveals a non-linear relationship between model size and performance [11]. The InternVL3 family exhibits a “Goldilocks effect,” [16] where the mid-sized 8B model excels in Accidents & Behaviors (0.83) and Natural Disasters (0.81), yet underperforms in Medical Emergencies (0.59) compared to both smaller 2B (0.75) and larger 14B (0.77) variants (see Table 4).
This suggests categories with obvious visual danger cues benefit from mid-sized models balancing perception and reasoning. Conversely, medical emergencies—requiring finer distinctions—perform better with either very small models (efficient pattern-matching) or large models (enhanced reasoning). The precision-recall tradeoff varies inconsistently with scaling. In Qwen2.5-VL, precision improves modestly with size (0.51→0.56→0.57), but recall fluctuates (1.00→0.88→1.00), challenging assumptions about model scaling in safety tasks.
Contextual Overinterpretation dominated across all parameter scales (89-91%). Even the best-performing model (InternVL3-8B) showed similar patterns (90.3% Contextual Overinterpretation), indicating a fundamental limitation in contextual reasoning that persists regardless of size. These findings suggest emergency recognition requires specialized architectures or fine-tuning approaches tailored to each category’s unique challenges.
Figure 5: Media-based danger misclassification. Left: Drive-in theater thunderstorm scene. Right: Flood poster viewed by pedestrian. Both cases show models failing to recognize representational contexts.
@@
4.6 Universal Misclassifications
Most striking is that 10 of our 100 safe images (10%) were misclassified by all 14 evaluated models, revealing common triggers for overreaction across architectures and parameter scales. These universally misclassified images represent the most extreme cases of the overreaction problem, with Visual Misinterpretation being the dominant factor.
The Visual Misinterpretation errors manifested in two key ways. First, models consistently failed to recognize representational contexts, as shown in Figure 5, where dangerous elements were portrayed in media rather than occurring in reality. Models misclassified scenes showing thunderstorms on drive-in theater screens or flood imagery on posters because they failed to perceive the critical visual cues that indicated these were representations. Second, as illustrated in Figure 6, models confused visually similar but contextually distinct scenarios, such as mistaking ketchup for blood or training mannequins for real people in danger.
@@ 5 Discussion
5.1 Balancing Safety and Accuracy
Our findings highlight a fundamental tension in visual emergency recognition: the trade-off between sensitivity to potential dangers and precision in distinguishing genuine emergencies from safe scenarios. While high sensitivity is desirable, the substantial false positive rates we observed (28-48%) could undermine practical utility in real-world applications. False alarms in emergency services lead to resource misallocation and ”alarm fatigue,” where users begin to distrust system warnings. Conversely, increasing precision must consider the severe consequences of missed emergencies. This optimization challenge varies across domains - medical triage might justifiably prioritize recall, while home monitoring systems require higher precision for user trust. The higher false positive rates in medical emergencies (38%) compared to natural disasters (27%) further suggest that different domains require customized approaches to balance sensitivity and specificity.
5.2 Implications for Model Development
Our findings highlight potential directions for improving VLMs’ emergency recognition capabilities. Enhancing contextual reasoning remains crucial, as models currently struggle to appropriately interpret contextual nuances. We observed that mid-sized models can sometimes achieve performance comparable to or even slightly better than larger models, indicating that simply increasing model size may not always yield significant improvements. However, given our relatively small sample size and modest performance differences observed between model families, further investigation with larger and more diverse datasets is needed. Additionally, category-specific optimizations and contrastive learning approaches could enhance models’ differentiation abilities.
5.3 Limitations
Our binary classification methodology inherently limits the exploration of nuanced severity gradations within emergency scenarios, which are crucial for realistic emergency assessments. Moreover, although synthetic images allow precise control over visual elements essential for diagnostic purposes, they may include subtle artifacts not representative of the complexities present in real-world photographs, potentially affecting generalization.
Our evaluation protocol simplifies the inherently complex nature of real-world emergency decision-making, and judgments provided by GPT-4o, despite their consistency, may not always fully align with expert evaluations. The VERI dataset, comprising 200 images structured into 100 contrastive pairs, intentionally emphasizes diagnostic depth over breadth. Consequently, this focused design inevitably poses limitations regarding broader generalizability, particularly across diverse cultural and geographical contexts.
Controlled laboratory evaluations using curated image pairs may not entirely predict model performance under realistic deployment conditions, characterized by imperfect inputs and variability. However, despite these acknowledged limitations, the consistently low performance scores observed in the evaluated models underscore the diagnostic value of this dataset. Future research should seek to mitigate these limitations through larger-scale datasets, more detailed classification schemas, and validations conducted under less controlled, more realistic scenarios.
@@ 6 Related Work
For VLMs to function safely in real-world applications, accurate risk assessment is essential. While our research addresses everyday risk assessment across multiple domains, previous work has primarily focused on domain-specific applications.
In autonomous driving, Zhang et al. [25] shows even larger models struggle with safety cognition, exhibiting limitations similar to our “overreaction problem.”
Fraser and Kiritchenko [9]’s PAIRS dataset uses parallel images that differ only in demographic attributes to evaluate social biases in VLMs. While it focuses on bias detection, it shares our approach of using AI-generated contrastive image pairs with controlled variations. Similarly, Tu et al. [20]’s benchmarks evaluates VLMs’ hallucination of non-existent objects, representing another critical safety concern complementary to our overreaction problem.
Previous research has documented VLMs’ deficiencies in commonsense reasoning, with models failing at simple tasks like identifying a lemon from “tastes sour” [22], achieving only $<42\%$ performance compared to human performance of $83\%$. The digital twin modeling field offers a complementary perspective, where Yang et al. [21] highlights how VLMs enable more flexible safety assessment through zero-shot learning. Recent evaluations show inconsistent scaling patterns [25], aligning with our findings on the persistent ”overreaction problem” across architectures and parameter scales.
@@ 7 Conclusion
We introduce VERI, a contrastive benchmark evaluating VLMs’ emergency recognition capabilities. Our findings reveal a systematic ”overreaction problem” with false positive rates of 28-48% across models, primarily stemming from contextual overinterpretation (88-93% of errors). This bias persists regardless of model scale, with mid-sized models sometimes outperforming larger variants. These results highlight the need for targeted improvements in contextual reasoning rather than simply scaling model size, suggesting future research in specialized architectures and contrastive learning approaches for safety-critical visual assessment.
Appendix A Emergency Response Evaluation Details
@@
A.1 GPT-4o Evaluation Methodology
To assess the quality of model-generated emergency responses (Q2), we used GPT-4o as an automatic evaluator. The prompt shown in Figure 7 was used to ensure consistency and transparency in scoring. This prompt provides the model with a description of the emergency situation, a gold-standard reference response, and the candidate response from the VLM. GPT-4o is then asked to assign a score between 0 and 1 based on specificity, factual accuracy, and alignment with expert protocols.
We additionally verified a random 10% subset of responses with two human annotators, achieving inter-rater agreement of $\kappa=0.83$, confirming that GPT-4o judgments were largely consistent with expert evaluations.
@@ A.1 GPT-4o Evaluation Methodology
A.2 Category-Specific Performance Data
@@
The InternVL3 family exhibits increasing consistency across emergency categories as model size grows, with the 14B variant yielding nearly identical scores across all three categories (maximum deviation: 0.004). This suggests that larger InternVL3 models generalize better across diverse emergency types. In contrast, the Mistral family shows stronger category-specific preferences even at large scales, with Pixtral-Large performing markedly better on AB (0.70) and ND (0.70) than on PME (0.63), indicating potential limitations in medical response generalization.
Smaller models ($<$8B) exhibit larger performance differences across categories compared to larger models ($>$20B), with exceptions. Performance on Medical Emergencies shows the most consistent improvement with model size across families, suggesting that medical intervention knowledge particularly benefits from increased capacity. These patterns reinforce our main finding that emergency response (Q2) performance is more stable across categories than risk identification (Q1), but also highlight family-specific trends, such as Mistral models’ preference for AB and ND, and InternVL3’s balanced scaling.
@@
Reference Response: Immediately move the child away from the hot kettle… If the child has already touched the hot surface, run the affected area under cool water for 10-15 minutes. Do not apply ice, butter, or ointments… For serious burns with blistering, seek medical attention promptly…
@@
Reference Response: Surfers should immediately exit the water as lightning is extremely dangerous when combined with salt water… Follow the 30/30 rule: if thunder is heard within 30 seconds of seeing lightning, the storm is dangerously close. Exit the water quickly and move far from the shoreline toward substantial shelter… Wait at least 30 minutes after the last lightning strike before returning to the water…
@@
”…individuals surfing in what appears to be stormy weather, with lightning present… This situation poses immediate risks… **Immediate Evacuation**: Surfers should immediately exit the water and seek shelter as quickly as possible to avoid lightning strikes. **Avoid Tall Structures or Water**: Moving to a flat, low-lying area away from water, tall structures, or trees is crucial to minimize risk…” Low-Quality Model Response (Score: 0.2):
@@
A.3 Emergency Response and Evaluation
@@
Appendix B Detailed Contextual Overinterpretation Example
Our main analysis revealed that Contextual Overinterpretation accounts for 88-93% of model misclassifications across evaluated VLMs. This section provides additional examples of this systematic error pattern, illustrating how models exaggerate risks in safe scenarios across different categories. As shown in Figure 9, models frequently misinterpret harmless activities, from card tricks and eating spaghetti to gardening, as dangerous situations that require intervention. These examples demonstrate how VLMs can correctly identify visual elements but consistently fail to assess their contextual safety implications, revealing a persistent ”better-safe-than-sorry” bias that manifests across different visual domains and model architectures.
@@ Appendix B Detailed Contextual Overinterpretation Example
Accidents & Behaviors 326 85.3% 14.7%
Natural Disasters 298 100.0% 0.0%
Personal Medical 237 85.7% 14.3%
Table 6: Distribution of error types by emergency category. CO: Contextual Overinterpretation, VM: Visual Misinterpretation. Note that the Natural Disasters category exhibits exclusively Contextual Overinterpretation errors.
@@
Appendix C Detailed Error Pattern Analysis
Our in-depth analysis of risk identification (Q1) errors revealed a remarkably consistent distribution across both models and categories. Regardless of architecture or parameter count (2B–124B), all evaluated models exhibited a strong bias toward Contextual Overinterpretation (CO), which accounted for 88–93% of false positives.
This trend held across model sizes: CO rates were uniformly high across scale groups—91.4% for 0–5B models, 90.2% for 5–10B, 89.5% for 10–20B, and 90.9% for models above 20B. Such consistency suggests that limitations in contextual reasoning are systemic within current VLM architectures and cannot be resolved by scaling alone. Even the top-performing model in terms of precision (InternVL3-8B) misclassified 90.3% of its false positives due to contextual overinterpretation, while Pixtral-Large (124B)—the largest model—had an even higher CO rate of 93.0%. Table 7 summarizes CO/VM distributions across all models.
Category-level analysis (Table 6) further supports this pattern. Natural Disasters exhibited exclusively CO errors (100%), indicating that models recognize elements like fire, smoke, or water but fail to reason about containment or safety context. Similar but slightly more diverse error profiles were observed in Accidents & Behaviors (85.3% CO) and Personal Medical Emergencies (85.7% CO), where CO errors still dominated but Visual Misinterpretations (VM) occasionally occurred.
Taken together, these findings suggest that while VLMs can detect visual features associated with danger, they struggle to weigh contextual cues accurately—particularly in ambiguous or representational scenarios.

### v3

@@ 3.8 Error Patterns and Analysis
Our analysis across 17 models revealed that Contextual Overinterpretation was dominant, accounting for 88-98% of all misclassifications regardless of model architecture or scale. Notably, in the Natural Disasters category, 100% of the errors were Contextual Overinterpretation, suggesting that models can correctly identify elements like fire or water but consistently fail to assess their contextual safety.
These findings suggest that current VLMs can detect potentially hazardous elements, but lack the nuanced reasoning to assess whether these elements pose actual dangers in specific contexts, a critical limitation for safety applications. Detailed error pattern analysis across model sizes and categories, along with additional examples of contextual overinterpretation, is provided in the supplementary material. The dominance of contextual overinterpretation likely reflects training data imbalance, where dangerous visual elements are predominantly paired with actual threats, combined with safety alignment methods that may inadvertently reinforce “better-safe-than-sorry” behaviors by penalizing false negatives more heavily than false positives.
@@ 3.9 Model Size and Category Effects
Interestingly, the InternVL3 8B model outperforms both smaller and larger variants in Accidents & Behaviors (0.825) and Natural Disasters (0.805), but shows the opposite pattern in Medical Emergencies (0.593 vs. 0.753 for 2B and 0.767 for 14B), suggesting category-specific scaling effects (see Table 4).
This suggests categories with obvious visual danger cues benefit from mid-sized models balancing perception and reasoning. Conversely, medical emergencies, which require finer distinctions, perform better with either very small models (efficient pattern-matching) or large models (enhanced reasoning). The precision–recall tradeoff varies inconsistently with scaling. In Qwen2.5-VL, precision improves modestly with size (0.510→0.554→0.589→0.652), but recall fluctuates (1.00→0.88→0.89→0.90), challenging assumptions about model scaling in safety tasks.
Contextual Overinterpretation dominated across all parameter scales (89–92%). Even the best-performing model (InternVL3-8B) showed similar patterns (90.3% Contextual Overinterpretation), indicating a fundamental limitation in contextual reasoning that persists regardless of size. Notably, commercial models exhibit even higher CO rates (90.7–98.1%), demonstrating that this systematic bias persists across different development approaches and scales. These findings suggest emergency recognition requires specialized architectures or fine-tuning approaches tailored to each category’s unique challenges.
3.10 Universal Misclassifications
Notably, 7 of our 100 safe images (7%) were misclassified by all 17 evaluated models, revealing common triggers for overreaction across architectures and parameter scales. These universally misclassified images represent the most extreme cases of the overreaction problem, with Visual Misinterpretation being the dominant factor. The Visual Misinterpretation errors manifested in two key ways. First, models consistently failed to recognize representational contexts, as shown in Figure 6, where dangerous elements were portrayed in media rather than occurring in reality. Models misclassified scenes showing thunderstorms on drive-in theater screens or flood imagery on posters because they failed to perceive the critical visual cues that indicated these were representations. Second, as illustrated in Figure 7, models confused visually similar but contextually distinct scenarios, such as mistaking ketchup for blood or training mannequins for real people in danger.
Figure 6: Media-based danger misclassification. Left: Drive-in theater thunderstorm scene. Right: Flood poster viewed by a pedestrian. Both cases show models failing to recognize representational contexts.
@@ 4 Discussion
Balancing Safety and Accuracy Our findings reveal a fundamental tension in emergency recognition: the trade-off between sensitivity and precision. The substantial false positive rates (31-96%) could undermine practical utility through resource misallocation and alarm fatigue, yet reducing sensitivity risks missing genuine emergencies. This optimization challenge varies by domain—medical triage justifiably prioritizes recall, while home monitoring requires higher precision for user trust. Cross-category variation in false positive rates suggests domain-specific calibration is necessary.
Implications for Model Development The dominance of contextual overinterpretation indicates that enhancing contextual reasoning, rather than visual encoding, should be prioritized. Mid-sized models sometimes matching larger variants suggests that scaling alone is insufficient. Even conservative prompting, requiring “clear and immediate risk,” merely trades precision for recall without addressing fundamental bias in learned representations. Category-specific optimizations and contrastive learning could improve discrimination between visually similar scenarios.
Limitations Our binary classification methodology limits exploration of nuanced severity gradations. While synthetic images enable precise control for diagnostic evaluation, they may contain artifacts affecting generalization. Real-world validation (25 pairs) confirms pattern consistency but remains limited in scale. The evaluation protocol simplifies complex emergency decision-making, and automated GPT-4o judgments may not capture all expert nuances. Despite the VERI dataset’s limited size (250 images total), consistent patterns across all models suggest diagnostic value. Future work should pursue larger-scale validation across diverse cultural and geographical contexts.
@@ 5 Related Work
For VLMs to function safely in real-world applications, accurate risk assessment is essential. While our research addresses everyday risk assessment across multiple domains, previous work has primarily focused on domain-specific applications. In autonomous driving, Zhang et al. [27] shows even larger models struggle with safety cognition, exhibiting limitations similar to our “overreaction problem.” Fraser and Kiritchenko [11]’s PAIRS dataset uses parallel images that differ only in demographic attributes to evaluate social biases in VLMs. While it focuses on bias detection, it shares our approach of using AI-generated contrastive image pairs with controlled variations. Similarly, Tu et al. [21]’s benchmarks evaluates VLMs’ hallucination of non-existent objects, representing another safety concern complementary to our overreaction problem. Recent work by Lee et al. [15] found VLMs show greater vulnerability to real-world memes, revealing complementary safety concerns to our overreaction problem in contextual reasoning.
Previous research has documented VLMs’ deficiencies in commonsense reasoning, with models failing at simple tasks like identifying a lemon from “tastes sour” [23], achieving only $42\%$ performance compared to human performance of $83\%$. The digital twin modeling field offers a complementary perspective, where Yang et al. [22] highlights how VLMs enable more flexible safety assessment through zero-shot learning. Recent evaluations show inconsistent scaling patterns [27], aligning with our findings on the persistent “overreaction problem” across architectures and parameter scales.
@@ 6 Conclusion
We introduce VERI, a contrastive benchmark evaluating VLM emergency recognition through 100 synthetic and 25 real-world image pairs, and find a systematic overreaction problem, with high recall (70–100%) but low precision (false positives 31–96%), largely driven by contextual overinterpretation (88–98%) with several safe scenarios universally misclassified, indicating limits across scales. These findings challenge the assumption that scaling alone improves safety-critical performance. To prevent alarm fatigue and retain trust, future models must strengthen contextual reasoning and adopt training strategies that balance sensitivity and specificity for real-world deployment.
Appendix A Real-World Validation Details
Collection Methodology We systematically searched for real-world images matching our synthetic scenarios through web search, focusing on finding visually similar pairs that maintain the emergency-safe distinction. We based our search on the synthetic scenarios, attempting to find the closest real-world equivalents. However, finding pairs with identical contextual setup proved challenging, limiting our collection to 50 validation images (25 pairs).
Dataset Composition Due to the availability constraints of matching real-world scenarios, the distribution differs from our synthetic dataset:
  • Accidents & Unsafe Behaviors: 15 pairs (60%)
  • Natural Disasters: 9 pairs (36%)
  • Personal Medical Emergencies: 1 pair (4%)
The limited Medical Emergency pairs reflect the difficulty in finding real-world images that match our controlled synthetic scenarios while maintaining ethical standards and visual similarity.
Representative Examples Figure 8 presents example real-world pairs demonstrating the visual similarity maintained across categories.
Figure 8: Representative real-world contrastive pairs from each category.
Appendix B Emergency Response Evaluation Details
@@
B.1 GPT-4o Evaluation Methodology
Given the large volume of responses (1,700 total across 17 models), we employed GPT-4o as an automatic evaluator, following the widely-adopted ’LLM-as-a-judge’ methodology. The prompt shown in Figure 9 was used to ensure consistency and transparency in scoring. This prompt provides the model with a description of the emergency situation, a gold-standard reference response, and the candidate response from the VLM. GPT-4o is then asked to assign a score between 0 and 1 based on specificity, factual accuracy, and alignment with expert protocols. To validate this approach, we manually verified 400 responses (23.5% of total) with human annotators, achieving strong inter-rater agreement of $\kappa=0.77$.
@@ B.1 GPT-4o Evaluation Methodology
B.2 Category-Specific Performance Data
@@
The InternVL3 family exhibits increasing consistency across emergency categories as model size grows, with the 14B variant yielding nearly identical scores across all three categories (maximum deviation: 0.004). In contrast, the Mistral family shows stronger category-specific preferences, with Pixtral-Large performing markedly better on AB (0.700) and ND (0.697) than on PME (0.625).
Commercial models demonstrate exceptional performance, with Gemini-2.5-Flash achieving the highest scores across all categories (0.814-0.842), showing remarkably consistent high-quality responses. GPT-4o and Claude-4-Sonnet also outperform most open-source models, though with more variation across categories.
These patterns reinforce our main finding that emergency response (Q2) performance is more stable across categories than risk identification (Q1), but also highlight the superior consistency of advanced commercial models.
@@
Commercial Models
Gemini-2.5-Flash 0.842 0.814 0.830
GPT-4o 0.607 0.620 0.782
Claude-4-Sonnet 0.668 0.743 0.797
@@
Reference Response:
Immediately move the child away from the hot kettle… If the child has already touched the hot surface, run the affected area under cool water for 10-15 minutes. Do not apply ice, butter, or ointments… For serious burns with blistering, seek medical attention promptly…
@@
Reference Response:
Surfers should immediately exit the water as lightning is extremely dangerous when combined with salt water… Follow the 30/30 rule: if thunder is heard within 30 seconds of seeing lightning, the storm is dangerously close. Exit the water quickly and move far from the shoreline toward substantial shelter… Wait at least 30 minutes after the last lightning strike before returning to the water…
@@
”…individuals surfing in what appears to be stormy weather, with lightning present… This situation poses immediate risks… **Immediate Evacuation**: Surfers should immediately exit the water and seek shelter as quickly as possible to avoid lightning strikes. **Avoid Tall Structures or Water**: Moving to a flat, low-lying area away from water, tall structures, or trees is crucial to minimize risk…”
Low-Quality Model Response (Score: 0.2):
@@
Appendix C Cost-Sensitive Evaluation
To address reviewer concerns about cost-sensitive evaluation and calibration, we conducted a post-hoc threshold sweep over the predicted probabilities for Q1 (risk identification). Figure 11 reports the ROC and precision–recall curves for three representative models (Qwen2.5-VL-72B, InternVL3-8B, Phi-3.5-Vision). While Qwen2.5-VL-72B operates near random chance (ROC-AUC $\approx 0.50$, AP $\approx 0.50$), InternVL3-8B and Phi-3.5-Vision achieve substantially higher discrimination (ROC-AUC $\approx 0.72$$0.73$, AP $\approx 0.68$$0.76$). At the $F_{1}$-maximizing threshold, InternVL3-8B reaches $F_{1}\approx 0.70$ ($P\approx 0.78$, $R\approx 0.63$), and Phi-3.5-Vision reaches $F_{1}\approx 0.73$ ($P\approx 0.75$, $R\approx 0.71$). These results confirm that, even after optimal threshold selection, the overreaction problem persists: false positives remain dominant, indicating that better calibration alone is insufficient to solve this issue.
[…]
C.1 Emergency Response and Evaluation
@@
Appendix D Detailed Contextual Overinterpretation Example
Our main analysis revealed that Contextual Overinterpretation accounts for 88-98% of model misclassifications across evaluated VLMs. This section provides additional examples of this systematic error pattern, illustrating how models exaggerate risks in safe scenarios across different categories. As shown in Figure 12, models frequently misinterpret harmless activities, from card tricks and eating spaghetti to gardening, as dangerous situations that require intervention. These examples demonstrate how VLMs can correctly identify visual elements but consistently fail to assess their contextual safety implications, revealing a persistent ”better-safe-than-sorry” bias that manifests across different visual domains and model architectures.
@@ Appendix D Detailed Contextual Overinterpretation Example
Accidents & Behaviors 393 86.0% 14.0%
Natural Disasters 362 100.0% 0.0%
Personal Medical 280 86.8% 13.2%
Table 7: Distribution of error types by emergency category across all 17 evaluated models. CO: Contextual Overinterpretation, VM: Visual Misinterpretation. Note that the Natural Disasters category exhibits exclusively Contextual Overinterpretation errors.
@@
Commercial Models
Gemini-2.5-Flash 54 98.1% 1.9%
GPT-4o 54 90.7% 9.3%
Claude-4-Sonnet 66 93.9% 6.1%
@@
Appendix E Detailed Error Pattern Analysis
Our in-depth analysis of risk identification (Q1) errors revealed a remarkably consistent distribution across models, architectures, and development approaches. Regardless of whether models were open-source or commercial, and spanning parameter counts from 2B to 124B, all 17 evaluated models exhibited a pronounced bias toward Contextual Overinterpretation (CO), which accounted for 88.4–98.1% of false positives.
This trend held across model sizes: CO rates were uniformly high across scale groups—90.1% for 0–5B models, 90.0% for 5–10B, 89.3% for 10–20B, and 91.6% for models above 20B. Commercial models showed even higher CO rates, averaging 94.2% with individual rates ranging from 90.7% (GPT-4o) to 98.1% (Gemini-2.5-Flash). Such consistency suggests that limitations in contextual reasoning are systemic within current VLM architectures and cannot be resolved by scaling alone or through sophisticated commercial training approaches.
Even the top-performing model in terms of precision (InternVL3-8B) misclassified 90.3% of its false positives due to contextual overinterpretation, while Pixtral-Large (124B)—the largest model—had an even higher CO rate of 93.0%. Notably, Gemini-2.5-Flash exhibited the highest CO rate at 98.1%, suggesting that even advanced commercial models with extensive training struggle with contextual reasoning. Table 8 summarizes CO/VM distributions across all models.
Category-level analysis (Table 7) further supports this pattern. Natural Disasters exhibited exclusively CO errors (100%), indicating that models recognize elements like fire, smoke, or water but fail to reason about containment or safety context. Similar but slightly more diverse error profiles were observed in Accidents & Behaviors (86.0% CO) and Personal Medical Emergencies (86.8% CO), where CO errors still dominated but Visual Misinterpretations (VM) occasionally occurred.
Taken together, these findings suggest that while VLMs can detect visual features associated with danger, they struggle to weigh contextual cues accurately—particularly in ambiguous or representational scenarios. This limitation persists across both open-source and commercial systems, indicating a fundamental challenge in current VLM architectures.